# Rui Lia,  Anyao Wangb   Mingqing Zhaic

### v2

@@ 5 Concluding remarks
We may also consider a generalization of the Hereditarily Bounded Property, where the hereditary bound on the number of edges is not restricted to linear constraints. In his thesis, Guiduli posed the following problem:
@@
If for a graph $G$, $|E(H)|\leq c\cdot|V(H)|^{2}$ holds for all $H\subseteq G$, does it follow that $\rho(G)\leq 2c\cdot|V(G)|$? What can be said if the exponent of $2$ is replaced by some other constant less than $2$?
This issue has also been highlighted by Liu and Ning in their survey of unsolved problems in spectral graph theory [16]. It is worth noting that $\rho(G)\leq\sqrt{2c}n$ serves as a trivial upper bound, and that $2c$ would be best possible, as seen by Wilf’s theorem. If this question were true, then the spectral Erdős-Stone-Simonovits theorem [17] would follow from the Erdős-Stone-Simonovits theorem [9, 10] and Wilf’s theorem would follow from Turán’s theorem [19].

### v3

@@ 5 Concluding remarks
We may also consider a generalization of the Hereditarily Bounded Property, where the hereditary bound on the number of edges is not restricted to linear constraints. In his thesis, Guiduli posed the following problem.
@@
Given a graph $G$, if $|E(H)|\leq c\cdot|V(H)|^{2}$ holds for all $H\subseteq G$, does it imply that $\rho(G)\leq 2c\cdot|V(G)|$? What can be said if the exponent of $2$ is replaced by some other constant less than $2$?
This issue has also been highlighted by Liu and Ning in their survey of unsolved problems in spectral graph theory [16]. It is worth noting that $\rho(G)\leq\sqrt{2c}n$ serves as a trivial upper bound, and that $2c$ would be best possible, as seen by Wilf’s inequality [20], i.e., $\rho(G)\leq(1-\frac{1}{\omega})n$, where $\omega$ denotes the clique number of $G$. If this question were true, then the spectral Erdős-Stone-Simonovits theorem [17] would follow from the Erdős-Stone-Simonovits theorem [9, 10] and Wilf’s theorem would follow from Turán’s theorem [19].